# When is Multicalibration Post-Processing Necessary?

### v1

@@ 3 Experiments on Tabular Datasets
We begin our investigation with tabular data. All our results are computed with a mean and standard deviation over 5 train / validation splits. Although simpler than vision or language data, tabular data is an important and realistic setting which many algorithmic fairness practitioners encounter throughout the health, criminal justice, finance, and other sectors (Barda et al., , 2021; Obermeyer et al., , 2019; Barenstein, , 2019). As our base predictors in this setting, we consider multilayer perceptron NNs (MLPs), DecisionTree and RandomForest, SVMs, NaiveBayes, and Logistic Regression. We defer dataset, group, and model details to Appendix C and Section E.1, but note that our datasets span from 10K examples (MEPS) to 200K (ACSIncome). We instill the following insights from running multicalibration post-processing algorithms over 40,000 times on more than 1,000 separately trained models.
@@
In Figure 1, we show the performance of every choice of multicalibration algorithm (corresponding to each choice of aforementioned hyperparameters) for MLPs on two datasets. We find that ERM performs nearly as well as the best set of hyperparameters across our wide parameter sweep. This is seen broadly across all of our tabular datasets for models which one may expect to be calibrated in practice, such as logistic regression or random forests [^2] We use a RF which predicts a probability corresponding to the fraction of positive points at the leaf.. We include the complete plots of all multicalibration runs versus ERM in Section F.1.
We numerically verify Observation 1 by inspection of Figure 2. This table corresponds to the best choice of hyperparameters (according to maximum group-wise smECE on a validation dataset) of each method tested on the CreditDefault dataset. We find that HKRR and HJZ show no improvements to max ECE or max smECE for MLPs, random forests, and logistic regression. Multicalibration gains are also marginal (0 to 0.01) on the Bank Marketing, ACSIncome, and Credit Default datasets (see Section F.2). On HMDA, however, multicalibration does seem to provide a noticeable improvement on the order of 0.03-0.05 for these models (Figure 21). We believe this is because ERM achieves worse calibration error on HMDA, possibly due to the increased difficulty of the dataset.
@@
Across our datasets, we find that SVMs, Decision Trees, and Naive Bayes almost always have their max smECE error improve by 0.05 or more using post-processing. We also point out the relatively strong performance of isotonic regression and other traditional calibration methods across datasets and models. For example, isotonic regression provides nearly all the improvements (up to 0.01 error) of the multicalibration algorithms when applied to Naive Bayes in Figure 2. On the Credit default dataset in Figure 22, isotonic regression is — when considering standard deviation — tied with the optimal multicalibration post-processing algorithms for SVM and Naive Bayes. We have similar findings for the MEPS dataset and Random Forests trained on the HMDA dataset. Platt scaling and isotonic regression are desirable because they are parameter-free methods working out of the box without tuning, are simple for practitioners to implement, and further do not require large parameter sweeps to find effective models.
@@
Figure 2: Best performing HKRR and HJZ post-processing algorithm hyperparameters (selected based on validation max smECE) compared to ERM on the MEPS dataset. Calibrated models (MLP, RandomForest, LogisticRegression) need not be post-processed to achieve multicalibration. However, uncalibrated models (SVM, Decision Trees, NaiveBayes) do benefit from multicalibration post-processing algorithms.
Observation 3: A practitioner utilizing multicalibration post-processing may have to trade off worst group calibration error and overall accuracy.
Due to the necessity of using hold-out data for running multicalibration post-processing, practitioners may have to choose between accuracy and worst group calibration error. For example, in Figure 3 we show that running multicalibration post-processing for MLPs on the HMDA dataset has a different optimal calibration fraction when considering accuracy and worst group calibration error as separate objectives. In fact, in this example, improving multicalibration error comes at a cost to accuracy of about 2%. Although small, this does indicate that the decision to use multicalibration post-processing here should be context dependent. Additional examples of this tradeoff include Decision Trees or Logistic Regression on most datasets (Figures 25 and 27). Plots for each dataset and model are in Section F.3.
Refer to captionRefer to caption
Figure 3: Hold-out calibration fraction vs. worst group calibration error (left) and accuracy (right) for MLPs on HMDA. Lowering worst group calibration error may come at a cost of model accuracy.
@@
To illustrate an example, ERM for Decision Trees on MEPS yields a worst group calibration error of 0.261 if calibration is measured using ECE, and 0.166 if using smECE (see Figure 2). This has important consequences for practitioners: if selecting the best model based on only the worst subgroup calibration error, the choice of calibration metric used will impact the choice of model. If selecting a post-processing algorithm based on ECE, HJZ may seem like a reasonable choice since it has a worst group ECE calibration error of 0.156. However, when using smECE to measure performance, HJZ does not significantly improve upon ERM.
We initially posited that this is because of the smaller sample sizes in some tabular datasets, but we note that even on the ACSIncome dataset with 200K examples, we find a significant difference of $0.1$ between measuring the overall calibration error of SVMs with ECE vs. smECE (Figure 19). More generally, to avoid issues stemming from ECE bin choice, we recommend that practitioners utilize the smECE calibration measurement tool [^3] https://github.com/apple/ml-calibration due to its theoretical guarantees and stability across our experiments.
Observation 5: There is no clearly dominant algorithm between HKRR and HJZ on tabular data, but HJZ is more robust towards the tested choice of hyperparameters.
Over the 30 experiments across all tabular datasets with all base models (Section F.2), HJZ and HKRR had statistically distinguishable performance on 24 runs. Amongst these 24, HJZ performed better 7 times. Nonetheless, we observe that the algorithm of HJZ is usually less sensitive to hyperparameter changes. In Figure 1 for example, the green points corresponding to hyperparameter choices for HJZ are tightly concentrated around ERM. We observe similar phenomena throughout additional model and dataset plots in Section F.1. Practitioners wishing to apply smaller hyperparameter searches over multicalibration algorithms may consider HJZ a suitable option, even if it gives slightly suboptimal worst group calibration error.
@@ 4 Experiments on Language and Vision Datasets
In this section, we evaluate the ability of multicalibration post-processing to improve upon the multicalibration of vision transformers, DistilBERT, ResNets, and DenseNets on a collection of image and language tasks. Our goal is to understand if multicalibration post-processing can help in more complicated, large model regimes within both the train-from-scratch and pre-trained paradigms.
@@
In all six of our experiments on image / language data (Section G.2), HKRR either matched or significantly beat the performance of HJZ. Note that we use the same parameter sweeps for HKRR and HJZ over image/language datasets that we used for the tabular datasets (see Appendix D), and leave open the possibility that HJZ may require a larger hyperparameter sweep to achieve good performance on more complex data.
Observation 7: Multicalibration post-processing can improve worst group smECE relative to an ERM baseline by 50% or more.
HKRR improved worst group calibration error in five out of six experiments. Among these five, the least improvement we saw was HKRR decreasing the worst group smECE of ERM from 0.06 to 0.043 (DistilBERT on Civil Comments). The greatest improvement we saw was from 0.07 to 0.02 (ViT on Camelyon17) and 0.09 to 0.05 (ResNet-56 on Amazon Polarity). These examples all appear in Figure 4. A full collection of tables and plots can be found in Section G.2.
@@
Among nearly all of our experiments on larger datasets of more than 100k examples, we were not able to find any datasets where the metric used to measure calibration error would change the outcome of chosen model (c.f. Observation 4).
Refer to captionRefer to caption
@@
Figure 4: Top: Test accuracy vs. maximum group-wise calibration error (smECE) averaged over three train/validation splits for ViT and DenseNet on Camelyon17. Multicalibration post-processing has scope for improvement in both settings, and does so with limited loss in accuracy. Bottom: Impact of multicalibration post-processing algorithms for Civil Comments (DistilBERT) and Amazon Polarity (ResNet-56). Multicalibration post-processing and isotonic calibration both offer improvements to worst group calibration error.
5 Takeaways for Practitioners and Experimental Limitations
@@
This directly motivates our next takeaway: Current multicalibration post-processing algorithms—when applied to calibrated models like neural networks—are extremely sensitive towards choice of hyperparameters, since the potential “scope” of improvement is on the scale of 0.02 to 0.03 smECE. The optimal hyperparameter choice for each algorithm largely varies by dataset and base model, and it takes quite a bit of granular searching to find the best performing algorithm, or indeed, an algorithm which improves upon ERM at all. For example, the optimal HJZ algorithm used at least 15 different hyperparameter configurations across only our 30 tabular experiments (when considering calibration fraction as an additional parameter); HKRR has similar sensitivity issues. Further, many hyperparameter choices do not seem to improve upon the ERM base model—for example, see DenseNet-121 in Figure 4 or the full multicalibration post-processing plots in Section G.1—making a significant portion of the hyperparameter sweeps not useful to perform. Since training HJZ or HKRR on a holdout of 100K examples can take 1-2 hours, it can be several hours before a suitable choice of hyperparameters is found. This computational cost is exacerbated in the larger regimes where multicalibration may be most useful, which poses a major obstacle for practical applications of either HKRR or HJZ.
@@
As a direct stopgap measure, we recommend running and evaluating traditional calibration methods. As we point out in Observation 2, post-processing algorithms like isotonic regression can achieve nearly the performance of multicalibration on tabular data. Isotonic regression also directly improves worst group calibration error over ERM in 4/6 of our experiments on larger models (see, e.g., Figures 4, G.2). Due to the fact that it is efficient and parameter free, we do not see a downside to running Isotonic regression (or any other calibration method) and testing if the maximum group-wise calibration error is beneath a desired threshold.
5.1 Limitations and Conclusion

### v2

@@ 3 Experiments on Tabular Datasets
We begin our investigation with tabular data. Although simpler than vision or language data, tabular data is an important and realistic setting which many algorithmic fairness practitioners encounter throughout the health, criminal justice, and finance sectors (Barda et al., , 2021; Obermeyer et al., , 2019; Barenstein, , 2019). As our base predictors in this setting, we consider multilayer perceptron NNs (MLPs), decision trees and random forests, SVMs, naive Bayes, and logistic regression. We defer dataset and group details to Appendix C, and model details to Section E.1. We note here that our datasets span from 10K examples (MEPS) to 200K (ACS Income), and we vary the size of the calibration set between 5% to 100% of the available training data. All of our results are computed with a mean and standard deviation over 5 train / validation splits. We instill the following insights from running multicalibration post-processing algorithms over 40,000 times on more than 1,000 separately trained models.
@@
In Figure 1, we show the performance of every choice of multicalibration algorithm (corresponding to each choice of aforementioned hyperparameters) for MLPs on three datasets: MEPS, Credit Default, and ACS Income. We find that ERM performs nearly as well — in terms of worst group calibration error — as the best set of hyperparameters for HJZ and HKRR across our wide parameter sweep. This is seen broadly across all of our tabular datasets for models which one may expect to be calibrated in practice, such as logistic regression or random forests. [^4] We use the Scikit-learn random forest implementation, which predicts a probability corresponding to the fraction of positive points at the leaf. We include the complete plots of all multicalibration runs versus ERM in Section F.1.
We provide further evidence for Observation 1 by inspection of Figure 2. This table corresponds to the best choice of hyperparameters (according to maximum group-wise smECE on a validation dataset) of each method tested on the MEPS dataset. We find that HKRR and HJZ show no statistically significant improvements to max smECE for MLPs, random forests, and logistic regression. The gains offered by HKRR and HJZ in terms of worst group calibration error are also marginal (0 to 0.01) on the Bank Marketing, ACS Income, and Credit Default datasets (see Section F.2). On HMDA, however, multicalibration does seem to provide a noticeable improvement on the order of 0.03-0.07 for MLPs, random forests, and logistic regression (Figure 27). We believe this is because ERM achieves worse calibration error on HMDA, possibly due to the increased difficulty of the dataset.
@@
Across our datasets, we find that SVMs, decision trees, and naive Bayes almost always have their max smECE error improve by 0.05 or more using multicalibration post-processing. We also point out the relatively strong performance of isotonic regression and other traditional calibration methods across datasets and models. For example, isotonic regression provides nearly all the improvements (up to 0.01 error) of the multicalibration algorithms when applied to naive Bayes in Figure 2. On the Credit Default dataset in Figure 28, isotonic regression is — when considering standard deviation — tied with the optimal multicalibration post-processing algorithms for SVM and naive Bayes. We have similar findings for the MEPS dataset and random forests trained on the HMDA dataset. Platt scaling and isotonic regression are desirable because they are parameter-free methods which work out of the box without tuning, are simple for practitioners to implement, and further do not require large parameter sweeps to find effective models.
@@
Figure 2: Best performing HKRR and HJZ post-processing algorithm hyperparameters (selected based on validation max smECE) compared to ERM on the MEPS dataset. Calibrated models (MLP, random forest, logistic regression) need not be post-processed to achieve multicalibration. However, uncalibrated models (SVM, decision trees, naive Bayes) do benefit from multicalibration post-processing algorithms. Cells highlighted in blue show the importance of the choice of metric for selecting the best post-processing method for decision trees. Metric choice — worst group ECE vs. worst group smECE — can change which of ERM or HJZ is preferable.
Observation 3: A practitioner utilizing multicalibration post-processing can potentially face a trade-off between worst group calibration error and overall accuracy. This is most salient in high calibration fraction regimes (40-80%).
Due to the necessity of using hold-out data for running multicalibration post-processing, practitioners may have to choose between accuracy and worst group calibration error. For example, in Figure 3 we show that running multicalibration post-processing for MLPs on the HMDA dataset has a different optimal calibration fraction when considering accuracy and worst group calibration error as separate objectives. In fact, in this example, improving multicalibration error comes at a cost to accuracy of about 2%. Although small, this does indicate that the decision to use multicalibration post-processing here should be context-dependent. Additional examples of this tradeoff include decision trees or logistic regression on most datasets (Figures 31 and 33). We note, however, that these tradeoffs are more apparent in the higher calibration fraction regimes, where most of the training data is held out for multicalibration post-processing. This regime is potentially less relevant to practitioners, who usually reserve most of the available data for base model training. Plots for each dataset and model are in Section F.3.
Refer to captionRefer to captionRefer to caption
Figure 3: (Left/Middle): Hold-out calibration fraction vs. worst group calibration error (left) and accuracy (right) for MLPs on HMDA. The left plot shows that the best max group smECE is achieved at 60% of data used for base model training, and 40% of data for multicalibration post-processing. However, the right plot shows that the best accuracy is achieved by using 90% of data for base model training, and the remaining for multicalibration post-processing. This means that a practitioner may face a tradeoff between worst-group calibration error and accuracy. The impact of calibration fraction for each dataset is available in Section F.3. (Right): Gap between measured smECE and ECE on every group for every experiment. As sample size increases, the two metrics become very similar. However, some variability exists at lower sample sizes.
@@
To illustrate an example, if a practitioner were selecting the best post-processing method for Decision Trees from Figure 2 based on ECE (see table cells highlighted in blue), HJZ may seem like a reasonable choice since it has a worst group ECE calibration error of 0.156. However, when using worst group smECE to measure performance, HJZ does not significantly improve upon ERM. This has an important consequence: if selecting the best model based on only the worst subgroup calibration error, the choice of calibration metric used will impact the choice of model.
In the rightmost plot of Figure 3, we also show each group’s sample size vs. the gap between measuring the group calibration error with smECE vs using ECE (over all datasets and groups). We find that as the group sample size increases, the gap between the metrics generally shrinks, and Observation 4 becomes less relevant. We note, however, that even on the ACS Income dataset with 200K examples, we find a significant difference of $0.1$ between measuring the overall calibration error of SVMs with ECE vs. smECE (Figure 25). More generally, to avoid issues stemming from ECE bin choice, we recommend that practitioners utilize the smECE calibration measurement tool [^5] https://github.com/apple/ml-calibration due to its theoretical guarantees and stability across our experiments.
Observation 5: When considering statistical significance, there is no clearly dominant algorithm between HKRR and HJZ on tabular data. However, HJZ is more robust towards the tested choice of hyperparameters. This may allow practitioners utilizing HJZ to find good solutions faster than using HKRR when post-processing simpler models such as naive Bayes or decision trees.
Over all tabular datasets and all base models (Section F.2), HJZ and HKRR had statistically distinguishable performance on 24 out of 30 cases. Among these 24 cases, HJZ performed better 7 times. Nonetheless, we observe that the HJZ family of algorithms is usually less sensitive to hyperparameter changes. In Figure 1 for example, the green points corresponding to hyperparameter choices for HJZ are tightly concentrated around ERM. We observe similar phenomena throughout additional model and dataset plots in Section F.1. Practitioners wishing to apply smaller hyperparameter searches over multicalibration algorithms may consider HJZ a suitable option, even if it gives slightly suboptimal worst group calibration error.
Additional Experiments. To understand how sensitive our observations are to our particular choice of group collection $G$, we also validate each of these observations with a new set of defined groups $G^{\prime}$, whose definitions are found in Section C.5. The full tabular results and plots for these new groups for each dataset is in Section G.1. Overall, the takeaway for most models (including MLPs) largely remains the same: it is difficult to find instances where multicalibration helps in a statistically significant way over ERM or some form of simple calibration (Observations 1 and 2). Further, where multicalibration does help, this help sometimes comes at a cost to accuracy (Observation 3).
@@ 4 Experiments on Language and Vision Datasets
In this section, we evaluate the ability of multicalibration post-processing to improve upon the multicalibration of vision transformers, DistilBERT, ResNets, and DenseNets on a collection of image and language tasks. Our goal is to understand if multicalibration post-processing can help in more complicated, large-model regimes within both the train-from-scratch and pre-trained paradigms.
@@
In all six of our experiments on image / language data (Section H.2), HKRR either matched or significantly beat the performance of HJZ. Note that we use the same parameter sweeps for HKRR and HJZ over image/language datasets that we used for the tabular datasets (see Appendix D), and leave open the possibility that HJZ may require a larger hyperparameter sweep to achieve good performance on more complex data.
Observation 7: On language and vision data, multicalibration post-processing can improve worst group calibration error relative to neural network ERM baselines by 50% or more. This stands in contrast to multicalibration post-processing for neural network ERM on tabular data, where we found no statistically significant improvements.
Over all language and vision datasets, HKRR improved worst group calibration error in 5 out of 6 cases. Among these 5, the least improvement we saw was HKRR decreasing the worst group smECE of ERM from 0.06 to 0.043 (DistilBERT on Civil Comments). The greatest improvement we saw was from 0.07 to 0.02 (ViT on Camelyon17) and 0.09 to 0.05 (ResNet-56 on Amazon Polarity). These examples all appear in Figure 4. A full collection of tables and plots can be found in Section H.2.
@@
Among nearly all of our experiments on larger datasets of more than 100k examples, we were not able to find any datasets where the metric used to measure calibration error would change the outcome of chosen model. This suggests that larger sample sizes close observable gaps between calibration measures (c.f. Observation 4).
Refer to captionRefer to captionRefer to caption
@@
Figure 4: Top: Test accuracy vs. maximum group-wise calibration error (smECE) averaged over three train/validation splits for ViT and DenseNet on Camelyon17, and DistilBERT on CivilComments. Multicalibration post-processing has scope for improvement in each setting, and does so with nearly no loss in accuracy. Bottom: Impact of multicalibration post-processing algorithms for Civil Comments (DistilBERT) and Amazon Polarity (ResNet-56). Multicalibration post-processing and isotonic regression both offer improvements to worst group calibration error. Full results are available in Section H.1.
5 Simplifying Data Partitioning with Data Reuse
As discussed in Section 2.2, throughout our experiments on tabular, vision, and language data in Sections 3 and 4, we held out a portion of data from training solely for running multicalibration post-processing. This was motivated by two facts: (1) multicalibration may require fresh samples for theoretical statistical guarantees; and (2) if a model already has low worst group calibration error on a holdout set $S$, that set $S$ cannot be used for post-processing since there are no “group calibration violations” that either HJZ or HKRR can correct. Fact (1) holds for any base model (neural networks, decision trees, random forests, etc.), while (2) only holds for models which we believe may always achieve perfect training loss or calibration error, like neural networks.
We investigate these two facts experimentally by asking whether a practitioner may reuse a portion of the base model training data for multicalibration post-processing. Such data reuse could be very convenient to practitioners already saddled with hyperparameter optimizations. To test for this, we run ERM on all available training data, and run each post-processing method on this same set of data. Again searching over all post-processing hyperparameters (now except for calibration fraction), we provide test results corresponding to hyperparameters that achieved the best validataion max smECE. We provide full results for all tabular datasets in Section F.4.
Observation 9: Reusing model training data for multicalibration post-processing can sometimes be competitive with holding out data for post-processing. However, it can also come at a steep cost to worst group calibration error.
In some cases, reusing training data can marginally improve max smECE over the setup in which we do not reuse data; this is true, for example, for MLPs post-processed with HKRR on the left of Figure 5. Here, reusing training data improves upon using a calibration holdout by 0.025. In the vast majority of cases, however, reusing training data either (1) does not improve upon utilizing a holdout calibration set (for most models on ACSIncome or HMDA); or (2) significantly hurts (for MEPS, CreditDefault, or BankMarketing). For example, many base models are significantly hurt by data reuse for multicalibration post-processing on the CreditDefault dataset (see the right plot of Figure 5), having their post-processed performance drastically drop by 0.05-0.1 worst group smECE. These results demonstrate that practitioners utilizing multicalibration algorithms in practice may be required to optimize over the calibration fraction holdout size in order to achieve competitive empirical performance.
Figure 5: Impact of reusing all model training data for multicalibration post-processing on HMDA (Left) and CreditDefault (Right) as measured by worst group calibration error (max smECE). Results vary; for HMDA, post-processing with reused data essentially performs as well as post-processing by holding out data for all models except random forest postprocessed with HKRR. However, on CreditDefault, we find that data reuse can harm post-processing across the board. Plots for each dataset available in Section F.4.1.
6 Takeaways for Practitioners and Discussion
In this section, we first provide reasonable recommendations to practitioners wishing to apply multicalibration algorithms in practice. In Section 6.1, we also discuss additional details on the subgroup selection problem, which practitioners applying post-processing methods may find helpful.
@@ 6 Takeaways for Practitioners and Discussion
This directly motivates our next takeaway: Current multicalibration post-processing algorithms—when applied to calibrated models like neural networks—are extremely sensitive towards choice of hyperparameters, since the potential “scope” of improvement is on the scale of 0.02 to 0.03 smECE. The optimal hyperparameter choice for each algorithm largely varies by dataset and base model, and it takes quite a bit of granular searching to find the best performing algorithm, or indeed, an algorithm which improves upon ERM at all. For example, the optimal HJZ algorithm used at least 15 different hyperparameter configurations across only our 30 tabular experiments (when considering calibration fraction as an additional parameter); HKRR has similar sensitivity issues. Further, many hyperparameter choices do not seem to improve upon the ERM base model—for example, see DenseNet-121 in Figure 4 or the full multicalibration post-processing plots in Section H.1—making a significant portion of the hyperparameter sweeps not useful to perform. Since training HJZ or HKRR on a holdout of 100K examples can take 1-2 hours, it can be several hours before a suitable choice of hyperparameters is found. This computational cost is exacerbated in the larger regimes where multicalibration may be most useful, which poses a major obstacle for practical applications of either HKRR or HJZ.
@@
As a direct stopgap measure, we recommend running and evaluating traditional calibration methods. As we point out in Observation 2, post-processing algorithms like isotonic regression can achieve nearly the performance of multicalibration algorithms on tabular data. Isotonic regression also directly improves worst group calibration error over ERM in 4/6 of our experiments on larger models (see, e.g., Figures 4, H.2). Due to the fact that it is efficient and parameter free, we do not see a downside to running Isotonic regression (or any other calibration method) and testing if the maximum group-wise calibration error is beneath a desired threshold.
6.1 Subgroup Design Considerations
Subgroup selection and design is inhererently important to running multicalibration post-processing, given that choice of subgroup collection can drastically impact both the post-processed model and resulting calibration metrics. For practitioners, there are (at least) two important properties of groups to consider during the group selection phase: minimum group size and the richness of the group collection. The minimum group size $\gamma$ is a parameter which has implications for the overall sample complexity of multicalibration. In particular, it introduces a $1/\gamma$ factor into known sample complexity upper and lower bounds (Hébert-Johnson et al., , 2018; Shabat et al., , 2020). Note that $\gamma\in[0,1]$ is the size — as a fraction of the dataset under a distribution $\mathcal{D}$ — of the smallest group in the collection $G$. Therefore, there is a tradeoff between the size of the smallest group considered, and the number of samples needed for good multicalibration generalization.
In our experiments, we restricted to groups which were >0.5% of the entire dataset ($\gamma=0.005$). This was a reasonable “sweet spot” for us: Without enough samples from a particular group, known multicalibration algorithms are prone to overfitting the training set and not providing desirable generalization performance. On the other hand, if groups are not small enough, then the guarantees provided by running multicalibration algorithm may not be much better than standard calibration methods. Note that we consider collections of groups with sizes spanning from 0.5% all the way to 70-80% of the data (see Appendix C for detailed information). We deem this range reasonably sufficient to capture the varying sizes of groups that a practitioner may desire to protect in practice.
Group richness is also an important factor in group design. As discussed in Section 2.2, our results are only relevant to the setting where (1) we have well-defined groups that we seek to protect which are defined by simple features or conjunctions of features; and (2) we run multicalibration post-processing with those same groups. These two points naturally give rise to a potentially promising direction: is it practically feasible to multicalibrate with respect to a richer class of subgroups in order to obtain better empirical performance with respect to the simpler groups which one may actually care about? For example, one could test whether multicalibrating with respect to the group of all halfspaces would provide improved worst group calibration error over the class of feature-defined groups than one may actually care about final performance for. [^6] Surprisingly, it is possible to multicalibrate with respect to the (rich) class of all halfspaces with access to an agnostic halfspace learner (Hébert-Johnson et al., , 2018). In general, such agnostic learners are computationally hard to obtain in theory, but usually easy to construct in practice (via ERM). We leave such exploration to future work.
7 Experimental Limitations and Conclusion